OpenStudy (anonymous):

the sum of two numbers is 16, and their difference is 2. write system of linear equations and solve the resulting system with x the larger number and y the smaller,

OpenStudy (anonymous):

I know what the first equation x+y=20 and x-y = 4. with the larger number being 20 and smaller 4. Is this correct

OpenStudy (anonymous):

hmm explain to me where the 20 and the 4 comes from

OpenStudy (anonymous):

the problem says let x be the larger number and y the smaller number

OpenStudy (anonymous):

"The sum of two numbers is 16" \[x+y=16\] "their difference is 2" \[x-y=2\]

OpenStudy (anonymous):

the sum of two number should be 20, and difference is 4, I had the numbers wrong

OpenStudy (anonymous):

ah excuse me, in this case you did perfect.

OpenStudy (radar):

Why are you saying the sum must be 20? The problem states the sum as 16.

OpenStudy (anonymous):

@radar she just corrected herself. "the sum of two number should be 20, and difference is 4, I had the numbers wrong"

OpenStudy (radar):

Oh @Spacelimbus, I see, you still deserve a medal, she can substitute the 20 for the 16 and 4 for the 2.

OpenStudy (anonymous):

Thanks a lot sir.